A commuter train carries 600 passengers each day from a town to a city. A one-way trip costs 100 per person. Market research reveals that 10 fewer people would ride the train with every 1 increase in the fare. What fare should be charged to get the largest possible revenue?

1
Expert's answer
2022-03-02T16:54:25-0500

Here the variables are

"x," the number of passengers,

"p," the price per ticket,

and "R," the revenue.

We have

"R = px"


We obtain


"R = p(600 -10(p-100)) ."

"R=-10p^2+1600p, p\ge100"

"a=-10<0, p_v=-\dfrac{b}{2a}=-\dfrac{1600}{2(-10)}=80"

"R(80)=-10(80)^2+1600(80)=64000"

The revenue has the absolute maximum with value of 64000 at price 80.